# Extension of the integrated HydroKinetic Model to BES RHIC and GSI-FAIR nuclear collision energies

### v2

@@ III.3 Bulk observables
This section presents our results using the tuned set of free parameters. The transverse momentum spectra for the lightest hadrons production in $0-5$% and $20-30$% centrality classes are shown in Fig. 2. It is evident that the model underestimates the $\bar{p}/p$ ratio, particularly noticeable in central collisions, and adjusting the free parameters does not fully resolve this discrepancy. This issue might be mitigated by considering dissipative terms in the baryon current (24), which could reduce the baryon chemical potential within the midrapidity region of the system. However, it could also indicate an overestimation of baryon stopping in the UrQMD model in combination with the hydrodynamization in the intermediate collision energy regime.
@@
Figure 4: iHKM results for the $m_{T}$ dependence of $R_{out}$, $R_{side}$ and $R_{long}$ for negatively charged pions in $5\%$ the most central collisions of gold nuclei at $\sqrt{s_{NN}}=14.5$ GeV. Left: iHKM results with crossover transition, and right: with first-order phase transition. Shaded regions cover values between STAR measurements for $11.5$ GeV and $19.6$ GeV [50]. Additionally, experimental results from the NA49 collaboration for Pb+Pb collisions at $17.3$ GeV [51] are included for reference.
@@
As one might see in Fig. 4, similarly to the elliptic flow $v_{2}$ results, iHKM performs much better if the equation of state with crossover-type transition is used. A large overestimate of $R_{long}$ in model calculations with Set 2 can be related to a relatively long hydrodynamic stage compared to the Set 1 scenario. This resulted from both by peculiarities of softer (lower pressure) EoS at low temperatures and later transition to the hadron stage due to the lower $\epsilon_{sw}$.
@@ III.4 Maximal emission times estimate
Figure 5: Two-pion and two-kaon $R_{\text{long}}(k_{T})$ dependence obtained from iHKM simulation with two different sets of parameters (see Table 1).
In both cases (Set 1 and Set 2), pions are emitted slightly earlier than kaons, consistent with our previous findings for ultrarelativistic energies. This observation can be attributed to the intense decay of $K^{*}(892)$ mesons during the afterburner stage.
Additionally, the fit indicates a stronger influence of collective flow on kaon emission compared to pions (smaller values of $\alpha$), leading to $k_{T}$ or $m_{T}$ scaling at large momenta.
@@ IV Summary
Summarizing, a model has been developed to describe the soft physics processes at the relativistic energies $2$ $50$ GeV per nucleon pair. A radical modification compared to the well-known iHKM model is the simulation of the initial stage of collisions at (relatively) low energy in the quasi-classical UrQMD model instead of the CGC+GLISSANDO representation for ultra-relativistic collisions. In addition to the basic theoretical foundations of the model, we also gave examples describing within its framework the spectra of pions, kaons, protons, and antiprotons for the intermediate energy of 14.5 GeV. Publications have also been prepared to describe the spectra, elliptical fluxes, and femtoscopy radii of the mentioned particles in the energy region from 7.7 to 39 GeV/nuclear pair, describing the data and aiming to investigate the possible phase transition interval.
@@ Appendix A Nonequilibrium distribution function in Milne coordinates
which exactly matches the expression in Eq. (43). Thus, in the continuous limit, introducing the smoothing kernel (8) does not violate conservation laws. Numerically, for RHIC BES energies, we observed violations of total energy and baryon charge that usually do not exceed $0.5\%$, attributed to the discreteness of the lattice with spacing $\Delta x=\Delta y=0.3$ fm, $\Delta\eta=0.05$, and a Gaussian radius parameter $R$ in the range of 0.50.50.50.51.01.01.01.0 fm.
@@ Acknowledgements.
The authors thank O. Vityuk and V. Naboka for their initial scientific contributions to this topic. They also acknowledge N. Rathod and H. Zbroszczyk for their support and valuable input in the development of the model. M.A. gratefully acknowledges support from the Simons Foundation (Grant SFI-PD-Ukraine-00014578). The research of Yu. S. was partially funded by IDUB-POB projects granted by WUT (Excellence Initiative: Research University (ID-UB)).

### v3

@@ III.3 Bulk observables
This section presents our results using the tuned set of free parameters. The transverse momentum spectra for the lightest hadrons production in 050-5% and 203020-30% centrality classes are shown in Fig. 3. It is evident that the model underestimates the $\bar{p}/p$ ratio, particularly noticeable in central collisions, and adjusting the free parameters does not fully resolve this discrepancy. This issue might be mitigated by considering dissipative terms in the baryon current (24), which could reduce the baryon chemical potential within the midrapidity region of the system. However, it could also indicate an overestimation of baryon stopping in the UrQMD model in combination with the hydrodynamization in the intermediate collision energy regime.
@@
Figure 5: iHKM results for the $m_{T}$ dependence of $R_{out}$, $R_{side}$ and $R_{long}$ for negatively charged pions in $5\%$ the most central collisions of gold nuclei at $\sqrt{s_{NN}}=14.5$ GeV. Left: iHKM results obtained with crossover equation of state, and right: with first-order phase transition. Shaded regions cover values between STAR measurements for $11.5$ GeV and $19.6$ GeV [68]. Additionally, experimental results from the NA49 collaboration for Pb+Pb collisions at $17.3$ GeV [69] are included for reference.
@@
As shown in Fig. 5, similar to the elliptic flow $v_{2}$ results, iHKM performs significantly better when the equation of state (EoS) with a crossover-type transition is used. The substantial overestimation of $R_{\text{long}}$ in model calculations with Set 2 can be attributed to the relatively long hydrodynamic stage compared to the Set 1 scenario. This difference arises from both the characteristics of a softer (lower-pressure) EoS at low temperatures and a later transition to the hadronic stage due to the lower switching energy density $\epsilon_{\text{sw}}$.
A qualitatively similar result for the $R_{\text{long}}$ component using the same equations of state, including the hadronic EoS, was reported in Ref. [72], which agrees with findings. However, quantitatively, in calculations with the crossover EoS, we achieve a better description of the “long” and “out” components, but a worse description in the “side” direction. A detailed beam energy scan study of pion interferometry within our model is left for future work.
@@ III.4 Maximal emission times estimate
Figure 6: Transverse momentum dependence of the two-pion and two-kaon $R_{\text{long}}(k_{T})$ radii obtained from the iHKM simulation with two different parameter sets (see Table 1). The dashed lines show the fits using Eq. (38).
In both cases (Set 1 and Set 2), pions are emitted slightly earlier than kaons, consistent with our previous findings at ultrarelativistic energies [13]. This time difference arises primarily from the substantial decay of $K^{*}(892)$ mesons during the afterburner stage. Additionally, the fit indicates a stronger influence of collective flow on kaon emission compared to pions, as reflected by the smaller values of $\alpha$. This leads to $k_{T}$ or $m_{T}$ scaling at higher transverse momenta.
A direct analysis of the last scattering times (i.e., particle emission times) from the UrQMD afterburner confirms the consistency of these results with those reported in Ref. [76]. Specifically, the maximal emission times correspond to particles with intermediate to high transverse momenta ($0.5<p_{T}<1.5~\mathrm{GeV}/c$), which is the momentum range used in the fits (the region where the spectra usually exhibit exponential decay). As observed previously at LHC energies, these emission times—especially for pions—are slightly smaller than the end times of the particlization stage. For instance, in the case of the crossover equation of state (Set 1), the pion emission time is approximately $6.6~\mathrm{fm}/c$, while the particlization ends around $9~\mathrm{fm}/c$, as shown in Table 2 and Fig. 1.
@@ IV Summary
Summarizing, a model has been developed to describe the soft physics processes at the relativistic energies $2$ - $50$ GeV per nucleon pair. A radical modification compared to the well-known iHKM model is the simulation of the initial stage of collisions at (relatively) low energy in the quasi-classical UrQMD model instead of the CGC+GLISSANDO representation for ultra-relativistic collisions. In addition to the basic theoretical foundations of the model, we also gave examples describing within its framework the spectra of pions, kaons, protons, and antiprotons for the intermediate energy of 14.5 GeV. Publications have also been prepared to describe the spectra, elliptical fluxes, and femtoscopy radii of the mentioned particles in the energy region from 7.7 to 39 GeV/nuclear pair, describing the data and aiming to investigate the possible phase transition interval.
In this work, we developed a formalism that, in principle, can serve as a basis for the analysis of collisions at very low relativistic energies, around $2~\mathrm{GeV}$ per nucleon pair, as relevant for the GSI HADES and future FAIR CBM experiments. However, several modifications become necessary in this low beam energy regime. First, as the energy of the nuclei decreases and the nuclear overlap time increases, interactions between nucleons during the earliest stages of the collision must be taken into account. Such features are implemented in modern transport codes. Furthermore, one cannot expect significant thermalization or hydrodynamization of nuclear matter, and the conventional hydrodynamic freeze-out criterion likely loses its applicability. As discussed in this paper, this situation may require a more computationally expensive but physically motivated construction of the particlization hypersurface based on the full energy-momentum tensor. A detailed analysis and application of our model to collisions at a few $\mathrm{GeV}$ per nucleon pair is planned for a separate publication.
@@ Appendix A Nonequilibrium distribution function in Milne coordinates
which exactly matches the expression in Eq. (45). Thus, in the continuous limit, introducing the smoothing kernel (8) does not violate conservation laws. Numerically, for RHIC BES energies, we observed violations of total energy and baryon charge that usually do not exceed $0.5\%$, attributed to the discreteness of the lattice with spacing $\Delta x=\Delta y=0.3$ fm, $\Delta\eta=0.05$, and a Gaussian radius parameter $R$ in the range of 0.50.51.01.0 fm.
@@ Acknowledgements.
The authors thank O. Vityuk and V. Naboka for their initial scientific contributions to this topic. They also acknowledge N. Rathod and H. Zbroszczyk for their support and valuable input in the development of the model. M.A. gratefully acknowledges support from the Simons Foundation (Grant SFI-PD-Ukraine-00014578). The research of Yu. S. was partially funded by IDUB-POB projects granted by WUT (Excellence Initiative: Research University (ID-UB)). This work is part of a project proposal submitted to the National Science Centre (OPUS 29 call, currently under review).